Consider the numbered balls shown. How many 2-digit numbers can be formed using these numbers if the balls are selected one at a time with replacement? 2,6,3,5,1,7,9
Answer #1

Answer:

Given Numbers ate 2, 6, 3, 5, 1, 7, 9

There are 2 positions and each position are not allowed to have repeat of the element,

There are total of 7 balls that means 7 numbers from which we have to choose

Therefore,

= 7P1 * 6P1 = 7 * 6 = 42

The number of choices (n) decreases by 1 with each new position because the choices cannot be repeated.

So, there are 42 2-digit numbers can be formed